# LLMBind: A Unified Modality-Task Integration Framework

### v5

@@
Abstract.
In the multi-modal domain, the dependence of various models on specific input formats leads to user confusion and hinders progress. To address this challenge, we introduce LLMBind, a novel framework designed to unify a diverse array of multi-modal tasks. By harnessing a Large Language Model (LLM) via Mixture-of-Experts (MoE), LLMBind processes multi-modal inputs and generates task-specific tokens, enabling the invocation of corresponding models to accomplish tasks. This unique approach empowers LLMBind to interpret inputs and generate outputs across various modalities, including image, text, video, and audio. Furthermore, we have constructed an interaction dataset comprising 400k instructions, which unlocks the ability of LLMBind for interactive visual generation and editing tasks. Extensive experimentation demonstrates that LLMBind achieves very superior performance across diverse tasks and outperforms existing models in user evaluations conducted in real-world scenarios. Moreover, the adaptability of LLMBind allows for seamless integration with the latest models and extension to new modality tasks, highlighting its potential to serve as a unified AI agent for modeling universal modalities.111The source code will be made publicly available.
1. Introduction
Recently, there has been significant progress in the field of Artificial Intelligence, including the Large Language Models (LLMs) (Chiang et al., 2023; Touvron et al., 2023) for natural language processing tasks and the diffusion models (Nichol et al., 2022; Tang et al., 2023; Liu et al., 2023a) for high-quality modality generation. Nevertheless, these models are typically limited to few modalities and tasks, highlighting the necessity to develop Multi-Modal Language Models (MLLMs) to integrate and address a broader range of tasks.
Existing methods tend to solve limited tasks and do not explore the unification of more tasks within a shared LLM. As shown in Fig. 1, we first use two characteristics to evaluate the efficacy of a framework, the updateability which quantifies the ease of replacing the frozen task-specific model, and the expandability which indicates the ability of the framework to extend to other tasks. Then, we summarize the existing MLLMs into three categories: 1) Encode-side Adapter, such as MiniGPT-4 (Zhu et al., 2023a) and LLaVA (Liu et al., 2023c), typically utilize encode-side adapters to handle inputs beyond language and can only perform the visual question-answering task; 2) Decode-side Adapter (Wu et al., 2023; Ghosal et al., 2023; Koh et al., 2023), as shown in Fig. 1(b), incorporates decode-side adapters for various generation models, enhancing the adaptation of LLMs to complex multimodal tasks like generation. However, these methods lack updateability as they require training a new adapter when introducing a new generation model; 3) Text Prompt Agent (Xia et al., 2023; Brooks et al., 2023), as depicted in Fig. 1(c), can directly generate desired text prompts for specific modality tasks without additional training. However, these approaches can only perform tasks that only have text prompts as additional inputs, and cannot solve more complex visual tasks such as image segmentation.
In this work, to achieve both updateability and expandability, we introduce LLMBind, a unified framework to integrate models for various modality tasks. As depicted in Fig. 1(d), LLMBind is capable of processing two modalities: image and text. It processes these inputs through a visual encoder (Radford et al., 2021) and LLM (Chiang et al., 2023), ultimately producing task-specific text prompts or token embeddings. For tasks such as video generation (Ding et al., 2021b; Rombach et al., 2022a), image generation (Nichol et al., 2022; Tang et al., 2023), audio generation (Liu et al., 2023a; Yang et al., 2023b), and image editing (Brooks et al., 2023; Fu et al., 2023), we utilize text prompts as input to existing pre-trained generation models, significantly reducing the training complexity of our framework and enhancing its updateability. For modality tasks like image segmentation (Lai et al., 2023; Liu et al., 2023b), classification (Yang et al., 2023a; Naeem et al., 2023), and detection (Ge et al., 2021) that cannot be solely addressed with text prompts, we extract relevant semantic visual information from images using learnable task-specific token embeddings. Subsequently, these embeddings are utilized as an auxiliary input to the corresponding modality-task models.
To guide the LLM to generate task-specific text prompts and token embeddings, we expand the vocabulary of LLM with learnable task-specific tokens, which can be divided into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. By training the LLM with specially designed instruction data pairs and task-specific tokens, we can instruct the model to discriminate the task type from user input and generate the corresponding text prompt or token embedding for a given multimodal task. Moreover, to better unify different modality tasks within the shared LLM, we introduce a highly effective technique called Mixture-of-Experts (MoE), which has been used successfully in recent studies (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Riquelme et al., 2021; Ahmed et al., 2016; Fedus et al., 2022). By incorporating this technique, our LLMBind can handle various modality tasks by utilizing different experts to collaboratively process multi-modal tasks, as shown in Fig. 2.
In practical applications, users often require multi-turn interactive generation and editing to achieve precise and satisfactory results. For example, in image editing, users need to apply various effects repeatedly to achieve their desired visual outcome. To this end, we develop a dataset comprising 400k instructional samples for interactive generation and editing tasks. Each sample, generated by ChatGPT, is formatted as a multi-round dialogue. This dataset alleviates the scarcity of instruction data for generation and editing tasks, equipping LLMBind with robust interaction capabilities.
Extensive experiments have demonstrated LLMBind’s superior performance in tasks such as referring segmentation, video generation, image generation, and audio generation, surpassing numerous specialized and multitask models. Additionally, LLMBind has exhibited remarkable performance in human evaluations of interactive generation tasks, exceeding traditional generation methods in terms of both the quality of generated data and its alignment with text. Overall, LLMBind demonstrates promising results in advancing the development of human-like MLLM and AI agents. Our Contributions to this work are as follows:
We introduce LLMBind, a unified framework for various modality tasks, encompassing image segmentation, generation, editing, as well as video and audio generation. We employ the MoE technique to effectively handle diverse tasks.
We devise a task-specific token strategy to seamlessly integrate different tasks into a unified MLLM, ensuring high model’s updateability and expandability for emerging tasks.
We curate an interactive dataset comprising 400k instructional data, enabling LLMBind to tackle multi-turn interactive visual generation and editing tasks.
2. Related Work
2.1. Cross-Modal Generation and Editing
@@
Multi-modal generation has remained a highly active research area in recent years. These models (Ramesh et al., 2022; Rombach et al., 2022b; Hong et al., 2022a; Kong et al., 2020; Yang et al., 2023c) allow the generation of corresponding images, videos, and audio based on specified text prompts. However, this approach requires users to directly write text prompts for image generation, lacking interactive capabilities. Another approach to cross-modal generation (Wu et al., 2023; Zeqiang et al., 2023), combines dialogue with generation, making it more user-friendly and aligned with current user needs. Given that generated images may not always meet user requirements, image editing has emerged as another active research area within the field of artificial intelligence-generated content. Representative works in this area include Text2live (Bar-Tal et al., 2022), Imagic (Kawar et al., 2023), Sdedit (Meng et al., 2021).
2.2. Cross-Modal Segmentation
@@
With the introduction of frameworks such as LanguageBind (Zhu et al., 2023b) and ImageBind (Girdhar et al., 2023), which combine multi-modal data in a shared semantic space, several works focusing on cross-modal understanding (Liu et al., 2023c; Zhu et al., 2023a; Yin et al., 2023; Jin et al., 2023) have emerged. Along with these advancements, there has been a growing interest in text-guided image referring segmentation, a new approach that enables users to interact with images using user instruction. Unlike traditional semantic segmentation, which assigns class labels to pixels, referring segmentation (Lai et al., 2023; Liu et al., 2023b) aims to segment target objects based on complex textual instructions.
@@
MLLMs Text-Con Img-Com Img-Gen Img-Edit Vid-Gen Aud-Gen Img-Seg Img-Det LLaMA (Touvron et al., 2023) ✔ ✗ ✗ ✗ ✗ ✗ ✗ ✗ Vicuna (Chiang et al., 2023) ✔ ✗ ✗ ✗ ✗ ✗ ✗ ✗ Flaminggo (Alayrac et al., 2022) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ BLIP2 (Li et al., 2023a) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ mPLUG-Owl (Ye et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ DetGPT (Pi et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✔ LLaVA (Liu et al., 2023c) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ Mini-GPT4 (Zhu et al., 2023a) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ Audiogpt (Huang et al., 2023) ✔ ✗ ✗ ✗ ✗ ✔ ✗ ✗ Otter (Li et al., 2023b) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ LISA (Lai et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✔ ✗ GLaMM (Rasheed et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✔ ✗ NExTGPT (Wu et al., 2023) ✔ ✔ ✔ ✗ ✔ ✔ ✗ ✗ LLMBind (Ours) ✔ ✔ ✔ ✔ ✔ ✔ ✔ -
2.3. Mixture-of-Experts
@@
2.4. Multimodal Large Language Models
Recently, the emergence of Large Language Models (LLM) has ushered AI into a new era. With the open-sourcing of LLM models like LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), numerous research works have successfully explored methods to extend LLMs to other modalities, including video, audio, and images. This has led to the development of a variety of related MLLM models, as shown in Table 1. For instance, Flamingo (Alayrac et al., 2022) employs cross-attention structures for visual-text interaction, while DetGPT (Pi et al., 2023) connects multimodal LLMs with detectors, enabling object detection based on user instructions. BLIP-2 (Li et al., 2023a) encodes input images into queries using the Q-Former structure. mPLUG-Owl (Ye et al., 2023) and LLaVA (Liu et al., 2023c) encode images into image features using image encoders, which are then combined with text embeddings for LLM learning. LISA (Lai et al., 2023) achieves referring segmentation by training LLMs to output mask token embeddings in an end-to-end manner. NExTGPT (Wu et al., 2023) implements multimodal generation by aligning the output of the adapter, which is added after the LLM, with the output text features from the text encoder of various generation models (Liu et al., 2023a; Rombach et al., 2022a).
3. Methodology
@@
3.1. LLMBind
Overview. The framework of LLMBind is illustrated in Fig. 3. It primarily consists of a vision transformer (Radford et al., 2021) as the image encoder, a large language model called Vicuna (Chiang et al., 2023), a SAM-like mask decoder as well as image, audio, and video diffusion models (Rombach et al., 2022a; Tang et al., 2023; Singer et al., 2022b) for generation and editing tasks.
@@ 3.1. LLMBind
Task-Specific Tokens. Different multimodal tasks often require distinct network structures, making it difficult to consolidate them within a shared LLM. To overcome this challenge, we expand the vocabulary of the LLM with specially designed, learnable task-specific tokens. As shown in Fig. 4 , these tokens can be classified into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. The first category includes tokens similar to standard text tokens, such as <gen> and </gen> for image generation, <vid_cap> and </vid_cap> for video generation, and <edit> and </edit> for image editing. These tokens appear in pairs, and the content between them represents the intermediate text prompt required for the corresponding modality task. In contrast, the second category is imbued with dense semantic information, diverging from traditional text tokens. For example, <seg> contains rich visual information related to segmented objects in an image. After decoding by a mask decoder, this token can be converted back into masks. This category also encompasses tokens like <detect> and <cls>, designed for extracting spatial coordinates and classification data in image-based tasks. This dual-token framework not only enhances the adaptability of the LLM across various modalities but also ensures scalability and extensibility.
Modality Task Integration. We seamlessly integrate various modality-task models in our framework using task-specific text prompts and token embeddings generated by LLM. For understanding tasks, we directly output text. In the generation and editing scenarios, we feed text prompts into the corresponding task model. Specifically, we use Stable Diffusion (Rombach et al., 2022a) for image generation, CoDi (Tang et al., 2023) for audio generation, ModelScopeT2V (Wang et al., 2023) for video generation, Instruct-pix2pix (Brooks et al., 2023) for image editing, thus ensuring high-quality, modality-specific output. For complex tasks such as referring segmentation, we use a single task-specific token embedding to learn the semantic information of the image features with respect to the segmentation instructions, which is then processed by a dedicated segmentation decoder. This Integration method is adaptable to various task requirements and has good extensibility.
3.2. LoRA Mixture-of-Experts
@@ 3.3. Training Objective
| Method | refCOCO | refCOCO+ | refCOCOg | |||||
| val | testA | testB | val | testA | testB | val(U) | test(U) | |
| MCN (Luo et al., 2020) | 62.4 | 64.2 | 59.7 | 50.6 | 55.0 | 44.7 | 49.2 | 49.4 |
| VLT (Ding et al., 2021a) | 67.5 | 70.5 | 65.2 | 56.3 | 61.0 | 50.1 | 55.0 | 57.7 |
| CRIS (Wang et al., 2022) | 70.5 | 73.2 | 66.1 | 62.3 | 68.1 | 53.7 | 59.9 | 60.4 |
| LAVT (Yang et al., 2022) | 72.7 | 75.8 | 68.8 | 62.1 | 68.4 | 55.1 | 61.2 | 62.1 |
| ReLA (Liu et al., 2023b) | 73.8 | 76.5 | 70.2 | 66.0 | 71.0 | 57.7 | 65.0 | 66.0 |
| X-Decoder (Zou et al., 2023a) | - | - | - | - | - | - | 64.6 | - |
| SEEM (Zou et al., 2023b) | - | - | - | - | - | - | 65.7 | - |
| LISA (Lai et al., 2023) | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 |
| LLMBind (ours) | 76.9 | 78.5 | 73.2 | 67.8 | 71.9 | 60.8 | 69.8 | 70.8 |
| Method | FID () |
| Specialized Models | |
| CogVideo (Ding et al., 2021b) | 27.10 |
| GLIDE (Nichol et al., 2022) | 12.24 |
| CoDi (Tang et al., 2023) | 11.26 |
| Multitask Models | |
| NExT-GPT (Wu et al., 2023) | 11.28 |
| LLMBind-SD1.5 | 16.41 |
| LLMBind-SD2 | 14.29 |
| LLMBind-GALIP* | 10.38 |
| Method | FID () | CLIPSIM () |
| Specialized Models | ||
| CogVideo (Hong et al., 2022b) | 23.59 | 0.2631 |
| Latent-VDM (Rombach et al., 2022a) | 14.25 | 0.2756 |
| Latent-Shift (An et al., 2023) | 15.23 | 0.2773 |
| Make-A-Video (Singer et al., 2022a) | 13.17 | 0.3049 |
| Multitask Models | ||
| NExT-GPT (Wu et al., 2023) | 13.04 | 0.3085 |
| LLMBind-Show-1 | 13.08 | 0.3072 |
| LLMBind-ModelScopeT2V | 11.09 | 0.2930 |
| LLMBind-HiGen | 8.60 | 0.2947 |
| Method | FD () | IS () |
| Specialized Models | ||
| DiffSound (Yang et al., 2023b) | 47.68 | 4.01 |
| AudioLDM-S (Liu et al., 2023a) | 29.48 | 6.90 |
| AudioLDM-L (Liu et al., 2023a) | 23.31 | 8.13 |
| Multitask Models | ||
| NExT-GPT (Wu et al., 2023) | 23.58 | 8.35 |
| LLMBind-CoDi | 22.90 | 8.77 |
| LLMBind-Auffusion | 21.99 | 10.57 |
3.4. Dataset

### v6

@@
Abstract
Despite recent progress in Multi-Modal Large Language Models (MLLMs), it remains challenging to integrate diverse tasks ranging from pixel-level perception to high-fidelity generation. Existing approaches often suffer from either restricted task extensibility or severe performance degradation due to modality interference. n this paper, we present LLMBind, an extensible framework that unifies multimodal tasks through a dual-pathway mechanism: In-Situ semantic embeddings for localization-sensitive tasks like semantic segmentation and Ex-Situ task-prompts for generation across image, video, and audio modalities. Additionally, we employ a Mixture-of-Experts (MoE) architecture to route task-specific tokens, thereby achieving modality disentanglement and mitigating negative transfer. We also curate a 400k multi-turn interactive dataset focused on iterative visual refinement to enable human-like interaction. Extensive experiments demonstrate that LLMBind achieves excellent performance across multiple perception and generation benchmarks while maintaining superior expandability.
1 Introduction
Artificial Intelligence is transitioning from specialized models toward general-purpose multimodal agents. While Vision-Language Models (VLMs) (Lai et al., 2023; Rasheed et al., 2024) excel in fine-grained perception and diffusion models (Nichol et al., 2022; Rombach et al., 2022b) provide high-fidelity synthesis. integrating both capabilities into a single framework is difficult. This difficulty stems from a fundamental architectural mismatch: perception requires tightly grounded In-Situ representations for spatial precision, whereas scalable generation favors Ex-Situ modular designs for flexibility and extensibility. Consequently, current approaches often specialize in one domain rather than achieving effective integration.
As depicted in Figure 1 (a), existing perception models like RexSeek (Jiang et al., 2025), LISA (Lai et al., 2023), PixelLM (Ren et al., 2024), PSALM (Zhang et al., 2024b) and GLaMM (Rasheed et al., 2024), embed visual signals directly into the latent space of the LLM. These In-Situ latent embeddings support high-precision tasks like referring segmentation, but the models lack the ability for content generation. Generation paradigms further divide into two categories. In-Situ generation approaches, as depicted in Figure 1 (b), such as NExT-GPT (Wu et al., 2023b), RISE-T2V (Zhang et al., 2025b), Sta-v2a (Ren et al., 2025), GILL (Koh et al., 2023), DreamLLM (Dong et al., 2023), MiniGPT5 (Zheng et al., 2023) and SEED-X (Ge et al., 2024), align LLM embeddings with generative latent spaces to enable synthesis. However, However, this creates a rigid coupling that restricts flexibility and complicates the adoption of new generative backbones. Conversely, the pure Ex-Situ generation, as shown in Figure 1 (c), including systems like Omnigen (Xiao et al., 2025) Visual ChatGPT (Wu et al., 2023a), Creatilayout (Zhang et al., 2025a) AudioGPT (Huang et al., 2024), InternGPT (Liu et al., 2023d), InstructPix2Pix (Brooks et al., 2023) and HuggingGPT (Shen et al., 2023), which invoke external tools via textual prompts. This offers strong modularity but suffers from a textual bottleneck, where sparse tokens cannot convey dense spatial information for pixel-level perception. Furthermore, jointly optimizing dense perceptual and abstract generative objectives triggers task interference, which degrades overall performance. These limitations demonstrate that reconciling distinct architectural and objective choices within a single framework remains a fundamental difficulty.
To resolve these contradictions, we introduce LLMBind, as shown in Figure 1 (d), a hybrid framework designed to balance In-Situ perception with Ex-Situ generation. Our core philosophy is a strategic bifurcation of multimodal tasks based on their intrinsic requirements for perception and generation granularity. For perception tasks, we employ learnable Semantic-Embedding Tokens to extract fine-grained visual information directly from the LLM’s latent space. Conversely, for generation-heavy tasks, we utilize Task-Prompt Tokens to steer frozen, state-of-the-art foundation models. Such a decoupled architecture empowers LLMBind with seamless plug-and-play expandability for emerging generative models without sacrificing the precision required for dense perception. To resolve optimization conflicts in shared-backbone multi-task learning, we propose a Task-Specific MoE architecture. The system routes task-specific tokens to dedicated LoRA experts to ensure explicit parameter-level isolation for different modalities. Consequently, LLMBind achieves functional decoupling between tasks, which effectively preserves task-specific expertise while leveraging the collective reasoning power of the LLM.
Furthermore, static instruction following does not meet the requirements of real-world collaboration. To solve this, we create an instruction-generative interleaved dataset with 400k samples. The dataset uses multi-turn formats to simulate refinement tasks, such as progressive image editing and multi-step visual generation. This allows LLMBind to achieve continuous visual refinement for human-like interaction. The main contributions are as follows:
We propose LLMBind, a dual-pathway framework unifying perception and generation via In-Situ embedding and Ex-Situ prompting.
We introduce a Task-Specific MoE strategy to mitigate modality interference through parameter-level isolation during joint multi-task training.
We release a 400k instruction-generative interleaved dataset for multi-turn visual refinement, including image editing and multi-step generation.
LLMBind outperforms specialized models in segmentation and generation, with human evaluation confirming high alignment with user instructions.
2 Related Work
2.1 Cross-Modal Generation and Editing
@@
Multi-modal generation has remained a highly active research area in recent years. These models (Ramesh et al., 2022; Rombach et al., 2022a; Hong et al., 2022b; Kong et al., 2020; Yang et al., 2023a; Xiao et al., 2025; Zhang et al., 2025a) allow the generation of corresponding images, videos, and audio based on specified text prompts. However, this approach requires users to directly write text prompts for image generation, lacking interactive capabilities. Another approach to cross-modal generation (Wu et al., 2023b; Zeqiang et al., 2023), combines dialogue with generation, making it more user-friendly and aligned with current user needs. Given that generated images may not always meet user requirements, image editing has emerged as another active research area within the field of artificial intelligence-generated content. Representative works in this area include Text2live (Bar-Tal et al., 2022), Imagic (Kawar et al., 2023), Sdedit (Meng et al., 2021).
2.2 Cross-Modal Segmentation
@@
With the introduction of frameworks such as LanguageBind (Zhu et al., 2023a) and ImageBind (Girdhar et al., 2023), which combine multi-modal data in a shared semantic space, several works focusing on cross-modal understanding (Liu et al., 2023c; Zhu et al., 2023b; Yin et al., 2023; Jin et al., 2023) have emerged. Along with these advancements, there has been a growing interest in text-guided image referring segmentation, a new approach that enables users to interact with images using user instruction. Unlike traditional semantic segmentation, which assigns class labels to pixels, referring segmentation (Lai et al., 2023; Liu et al., 2023a; Zhang et al., 2024a) aims to segment target objects based on complex textual instructions.
@@
MLLMs Text-Con Img-Com Img-Gen Img-Edit Vid-Gen Aud-Gen Img-Seg Img-Det LLaMA (Touvron et al., 2023) ✔ ✗ ✗ ✗ ✗ ✗ ✗ ✗ Vicuna (Chiang et al., 2023) ✔ ✗ ✗ ✗ ✗ ✗ ✗ ✗ Flaminggo (Alayrac et al., 2022) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ BLIP2 (Li et al., 2023b) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ mPLUG-Owl (Ye et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ DetGPT (Pi et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✔ LLaVA (Liu et al., 2023c) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ Mini-GPT4 (Zhu et al., 2023b) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ Audiogpt (Huang et al., 2023) ✔ ✗ ✗ ✗ ✗ ✔ ✗ ✗ Otter (Li et al., 2023a) ✔ ✔ ✗ ✗ ✗ ✗ ✗ ✗ LISA (Lai et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✔ ✗ GLaMM (Rasheed et al., 2023) ✔ ✔ ✗ ✗ ✗ ✗ ✔ ✗ NExTGPT (Wu et al., 2023b) ✔ ✔ ✔ ✗ ✔ ✔ ✗ ✗ LLMBind (Ours) ✔ ✔ ✔ ✔ ✔ ✔ ✔ -
2.3 Mixture-of-Experts
@@
2.4 Multimodal Large Language Models
Recently, the emergence of Large Language Models (LLM) has ushered AI into a new era. With the open-sourcing of LLM models like LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), numerous research works have successfully explored methods to extend LLMs to other modalities, including video, audio, and images. This has led to the development of a variety of related MLLM models, as shown in Table 1. For instance, Flamingo (Alayrac et al., 2022) employs cross-attention structures for visual-text interaction, while DetGPT (Pi et al., 2023) connects multimodal LLMs with detectors, enabling object detection based on user instructions. BLIP-2 (Li et al., 2023b) encodes input images into queries using the Q-Former structure. mPLUG-Owl (Ye et al., 2023) and LLaVA (Liu et al., 2023c) encode images into image features using image encoders, which are then combined with text embeddings for LLM learning. LISA (Lai et al., 2023) achieves referring segmentation by training LLMs to output mask token embeddings in an end-to-end manner. NExTGPT (Wu et al., 2023b) implements multimodal generation by aligning the output of the adapter, which is added after the LLM, with the output text features from the text encoder of various generation models (Liu et al., 2023b; Rombach et al., 2022b).
3 Methodology
@@
3.1 LLMBind
Overview. The framework of LLMBind is illustrated in Figure 3. It primarily consists of a vision transformer (Radford et al., 2021) as the image encoder, a large language model called Vicuna (Chiang et al., 2023), a SAM-like mask decoder as well as image, audio, and video diffusion models (Rombach et al., 2022b; Tang et al., 2023; Singer et al., 2022b) for generation and editing tasks.
@@ 3.1 LLMBind
Task-Specific Tokens. Different multimodal tasks often require distinct network structures, making it difficult to consolidate them within a shared LLM. To overcome this challenge, we expand the vocabulary of the LLM with specially designed, learnable task-specific tokens. As shown in Figure 2 , these tokens can be classified into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. The first category includes tokens similar to standard text tokens, such as <gen> and </gen> for image generation, <vid_cap> and </vid_cap> for video generation, and <edit> and </edit> for image editing. These tokens appear in pairs, and the content between them represents the intermediate text prompt required for the corresponding modality task. In contrast, the second category is imbued with dense semantic information, diverging from traditional text tokens. For example, <seg> contains rich visual information related to segmented objects in an image. After decoding by a mask decoder, this token can be converted back into masks. This category also encompasses tokens like <detect> and <cls>, designed for extracting spatial coordinates and classification data in image-based tasks. This dual-token framework not only enhances the adaptability of the LLM across various modalities but also ensures scalability and extensibility.
Modality Task Integration. We seamlessly integrate various modality-task models in our framework using task-specific text prompts and token embeddings generated by LLM. For understanding tasks, we directly output text. In the generation and editing scenarios, we feed text prompts into the corresponding task model. Specifically, we use Stable Diffusion (Rombach et al., 2022b) for image generation, CoDi (Tang et al., 2023) for audio generation, ModelScopeT2V (Wang et al., 2023) for video generation, Instruct-pix2pix (Brooks et al., 2023) for image editing, thus ensuring high-quality, modality-specific output. For complex tasks such as referring segmentation, we use a single task-specific token embedding to learn the semantic information of the image features with respect to the segmentation instructions, which is then processed by a dedicated segmentation decoder. This Integration method is adaptable to various task requirements and has good extensibility.
3.2 LoRA Mixture-of-Experts
@@ 3.3 Training Objective
3.4 Dataset